1. Solve algebraically the equation 15sin2x°=10cosx° for 0≤x<360.

Answer by stanbon(75887)

Solve algebraically the equation
15sin(2x)°=10cos(x)° for 0≤x<360.
---
sin(2x)/cos(x) = 2/3
----
(2sin(x)(cos(x))/cos(x) = 2/3
----
2sin(x) = 2/3
---
sin(x) = (1/3)
---
x = sin^1(1/3) = 19.47 degrees or 180-19.47 = 160.53 degrees
=================
Cheers,
Stan H.